Pixel:

Your annual health insurance premium is $3000. Your business pays 65%. If you are paid weekly, what is the paycheck deduction?

Pixel:

.65 x 3000 = 1950 3000 - 1950 = 1050 1050 / 52 = 20.19

TheSmartOne:

You see this calculation? .65 x 3000 = 1950 3000 - 1950 = 1050 You don't have to do all that. They're paying 65% of it. So how much are you paying then? The rest of it. Which is? 100% - 65% = 35% So then you can just do .35 * 3000/52 And you just save more time and still arrive at the correct answer

TheSmartOne:

You don't have to do .65 times 3000 And remember that number And then subtract it from 3000 You can just quickly do .35 * 3000 And you'll get the same number

TheSmartOne:

But nonetheless, good job, it's correct